OpenStudy (anonymous):

Solve: 2a/3 + 1/2 = (6a-1)/6 I hit a mental block involving equations with fractions...somebody care to explain how to do this one?

OpenStudy (anonymous):

the least common multiple of the denominators is 6, so multiply everything by 6: 4a+3=6a-1 4=2a a=2

OpenStudy (anonymous):

the LCD is 6, so multiply both sides by 6 thus getting 4a+3=6a-1 6a-4a=3+1 2a=4 a=2

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

you are welcome.